OpenStudy (adryin):

A. Write the first 5 terms of the recursively defined sequence. B. Is the sequence arithmetic, geometric, or neither? How do you know? C. For what values of x, if any, does the sequence converge? D. If the sum of the sequence is 5, what is the value of x?

OpenStudy (adryin):

\[A{n} = x/4 A{n-1}\] , A of 1 = 6

OpenStudy (mathmale):

Your presentation of this problem is vague and requires analysis by potential helpers:\[A{n} = x/4 A{n-1}\]

OpenStudy (mathmale):

\[A{n} = x/4 A{n-1}\]

OpenStudy (mathmale):

If you're going to use Equation Editor anyway, please use its subscript feature:\[A _{n} = x/4 A _{n-1}\]

OpenStudy (mathmale):

Unfortunately, I still don't understand how you obtain \[A _{n}\]

OpenStudy (mathmale):

from \[A _{n-1}\]

OpenStudy (mathmale):

Where does that x come from? If you could take a screen shot of this problem and post that image here, you might get more responses. Alternatively, draw the problem in the Draw utility, below.

OpenStudy (adryin):

X is originally from the sequence